# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v1

@@ 6. Multi Kernel Forward
Long KV Split. Multi-stream execution alone cannot eliminate execution bubbles in all cases, because some CTAs may have KV lengths that are orders of magnitude larger than others. We therefore adopt a KV-dimension splitting strategy similar to (Pan et al., 2025). Specifically, we split any CTAs with KV length exceeding the mean KV length for all CTAs of the batch into equal parts to keep KV length under the mean value. This shortens the completion time of the last finishing CTAs and improves overall SM utilization of kernel execution.
@@ 7. Output Merge
We implement a lightweight merge kernel that uses online softmax (Dao et al., 2022) to combine partial results at per-query granularity. Each CTA produces three per-query and per-head intermediates: a max score, a log-sum-exp accumulator, and a partial value-weighted sum. The merge kernel loads these intermediates from global memory, reduces them with online softmax, normalizes the accumulated value-weighted sum, concatenates all heads, and writes the final query output back to global memory. The small global read/write overhead for these intermediates is accounted for in the pack scheduler’s overhead analysis when deriving the end-to-end packing scheme (5.1).
@@ 8. Evaluation
In this section, we present the implementation of PAT (§8.1) and the experimental setup (§8.2). We then conduct a set of experiments to answer three key questions [^7] We also provide an overhead analysis in Appendix B.:
@@ 8.1. Implementation of PAT
We implement PAT as a full attention kernel for the decode stage with about 3k lines of Cutlass/CuTe (NVIDIA, 2025f, e) and C++ code. The multi-tile kernel (§5.2), multi-stream forward (§6), and merge kernel (§7) are built with Cutlass/CuTe. The asynchronous pack scheduler (§5.1) and API wrappers are implemented in C++. To overlap the data transfers with computation, all data movement from the global memory to the shared memory uses the cp_async primitive, together with double buffering (NVIDIA, 2025b). We expose the kernel API to Python via pybind11, and integrate it into vLLM (Kwon et al., 2023) (v0.9.0) as an off-the-shelf plugin with about 1.2k lines of Python code. To enable PAT in vLLM, only an environment variable VLLM_ATTENTION_BACKEND=PAT is required.
@@ 8.2. Experiment Setup
Models and Testbed. For end-to-end online serving (§8.4), we use two representative open-source models: Qwen3-8B (Yang et al., 2025) and Llama3-8B (AI@Meta, 2024). All experiments run on the NVIDIA A100-80GB-SXM4 GPU. Hardware info is summarized in Table 1. Software environment is CUDA 12.4 and PyTorch 2.7.0.
Baselines. We compare against five attention implementations spanning query-centric and KV-centric designs: (1) FlashAttention  (Dao et al., 2022; Dao, 2023) (v2.6.2): query-centric; maps each query to a CTA with a fixed tile config $(64,128)$. (2) FlashInfer  (Ye et al., 2025) (v0.2.5): query-centric; improves SM load balance by dynamic CTA partitioning; decoding tile config $(16,128)$. (3) FastTree  (Pan et al., 2025): KV-centric; uses a compute-oriented cost model to pack and reduce repeated KV loads; two tile configs $(64,32)$ and $(16,32)$. (4) RelayAttention  (Zhu et al., 2024): KV-centric; packs single-level shared prefixes across requests into CTAs to cut redundant memory accesses and runs them with FlashAttention’s kernel. (5) RelayAttention++: our extension of RelayAttention to exploit vLLM-style KV-cache reuse; it stores shared KV blocks from non-first-level prefixes in the same physical space so redundant KV loads can benefit from L2 cache, further improving performance (§8.3).
@@
Kernel-Performance Workloads (§8.3). To compare kernel performance under identical batch and prefix structures, we construct synthetic decode batches as input, following (Pan et al., 2025). Each decode batch is specified by $B$ and $L$. $B$ defines the prefix-tree structure and the number of leaves (i.e., batch size). e.g., $B=[1,4,16]$ yields two shared-prefix levels with 1 and 4 nodes, and 16 leaves. $L$ gives KV lengths per level. e.g., $L=[128,256,1024]$ sets level-1 and level-2 shared prefixes to 128 and 256 tokens, with 1024 non-shared tokens per request. We vary $(B,L)$ combinations to reflect different shared-prefix structures and batch settings as in Figure 9. Besides, we choose four head configurations (#heads, #kv_heads) common in Llama (AI@Meta, 2024), Qwen (Yang et al., 2025), and Gemma (Team, 2025) models: $(64,8)$, $(32,8)$, $(16,8)$, and $(32,32)$. The head dimension and data type are set to the commonly used 128 and FP16.
@@
End-to-end workloads (§8.4 and §8.5). We evaluate PAT under online serving using vLLM (Kwon et al., 2023) (v0.9.0) on two real-world traces. (1) toolagent (Qin et al., 2025): tool/agent workloads with task-specific system prompts (overall cache hit rate 59%). (2) conversation: combines the Meta-AI system instruction (total lengths 2517/2522 tokens for Llama3/Qwen3 tokenizers) with burstgpt prompts, following prior work (Pan et al., 2025). We randomize language and country fields in the system instruction to create a three-level prefix whose lengths are 46, 348, 2123 (Llama3) or 45, 351, 2126 (Qwen3). We use only the first 30 minutes of each trace due to cost limit.
@@ 8.3. Kernel Performance
Figure 10. End-to-end performance of PAT and the baselines under two models and two traces. Note that (1) RelayAttention++ lacks support for multiple first-level prefixes, so its results on toolagent trace are unavailable; (2) We place the results on Llama-3-8B toolagent and Qwen3-8B conversation in Appendix A due to page limits, and they exhibit similar trends.:
Overall results. Figure 9 reports normalized kernel performance (metric $\text{latency}^{-1}$, normalized to PAT) across decode-batch configurations. For configurations with shared prefixes ( 118), PAT achieves $22.7\times$, $12.0\times$, $2.0\times$, $12.9\times$ and $6.5\times$ speedups over FlashAttention, FlashInfer, FastTree, RelayAttention and RelayAttention++, respectively, across four attention-head settings. These gains arise from three factors: (1) a prefix-aware packing scheduler that cuts redundant global KV accesses (§5.1); (2) a multi-tile kernel combined with an online tile-size selector that adapts $(m,n)$ per CTA to better use bandwidth and on-chip memory (§5.2); and (3) multi-stream forward execution and long-KV splitting that mitigate execution bubbles caused by multi-tile kernels and KV-length dynamics (§6). Together, these designs make PAT consistently more efficient than the baselines.
Compared with query-centric kernels. Against query-centric FlashAttention and FlashInfer, PAT reduces attention latency by $12.3\%$$95.6\%$ on configurations with prefixes ( 118). The gap widens with larger batch size or longer shared prefixes (e.g., 45 and 67) because the one-query-per-CTA design of query-centric kernels forces repeated global KV loads that grow more costly in those cases. Even without shared prefixes ( 19 and 20), PAT’s multi-tile kernel combined with multi-stream forward reduces execution bubbles and yields $1.4\%$$21.2\%$ lower latency.
Compared with KV-centric kernels. FastTree, which also targets multi-level shared prefixes, remains the strongest baseline but is still $5.7\%$$103\%$ slower than PAT for two reasons: (1) its compute-oriented packing cost model is ill-suited to memory-bound decode attention (§2.3 and §3.2); and (2) its double-tile approach launches two kernels serially as in Figure 12(b), introducing execution bubbles. By contrast, PAT uses a memory-oriented packing strategy plus multi-stream forward to avoid these inefficiencies. RelayAttention++ cuts latency by $67.4\%$ versus RelayAttention, confirming that L2 plus KV-cache reuse reduces redundant global loads, yet RelayAttention++ is still $2.0\times$ slower than PAT, showing that L2 cache alone cannot fully eliminate redundant KV loads.
@@ 8.4. End-to-End Comparison
Overall trends. We compare PAT with three baselines across two models (Llama-3-8B and Qwen3-8B) and two real-world traces (conversation and toolagent) as in §8.2. Figure 10 shows the trend of mean TTFT, TPOT, and request e2e latency as we vary the request rate. Specifically, as in subfigures b1 to b2, PAT reduces TPOT at the same request rate by 14.9–56.4% over RelayAttention++, 13.6–73.9% over FlashAttention, and 31.6–83.4% over FlashInfer. Furthermore, as shown by the first and the third rows of Figure 10, the TPOT improvement allows PAT to finish incoming requests faster at the same request rate. This yields 8.6–97.2%, 7.9–99.6%, and 47.7–99.7% lower TTFT than the three baselines, and similar reductions in the e2e latency.
@@ 8.5. Ablation Study
Figure 12. CTA execution pipeline on $SM_{0}-SM_{5}$ with two tile size configurations (collected by PTX (NVIDIA, 2023)), where white space represents execution bubbles. (a) Multi-stream execution pipeline. (b) Serial execution pipeline.:
To evaluate the contribution of each design in PAT, we build the following ablation baselines: (1) PAT-compute, which adopts the cost model from FastTree (Pan et al., 2025) for the packing scheduler; (2) PAT-naive, which simply packs each node in the tree-structured block table into a CTA; (3) PAT-fixed, which disables the multi-tile kernel of PAT and instead uses the fixed tile configuration $(64,128)$ as in FlashAttention (Shah et al., 2024); (4) PAT-serial, which disables the multi-stream forward mechanism of PAT and adopts the serial multi-kernel execution similar to FastTree (Pan et al., 2025). We use the same synthetic traces as in §8.3 and adopt the attention head configuration of Llama-3-8B to compare PAT with ablation baselines.
@@
Importance of multi-tile kernel. As shown in Figure 11(a), enforcing a fixed tile size (PAT-fixed) increases attention latency by 39% compared to PAT. In the ablation workload, CTAs exhibit query sizes from 1 to 64 and KV lengths from 32 to 4096, which makes one-size-fits-all kernels highly inefficient. Prior kernels (Shah et al., 2024; Ye et al., 2025; Zhu et al., 2024; Yao et al., 2024) adopt such fixed designs, leading to padding overhead and execution bubbles. In contrast, PAT’s multi-tile kernel enables efficient adaptation to diverse CTA configurations.
@@ 9. Discussion
Prospects and Limitations. PAT leverages cross-request prefix sharing to improve bandwidth utilization and cut redundant global memory accesses in memory-bound decode attention. Its effectiveness depends on three factors: (1) hardware compute-to-bandwidth ratio. As GPUs become more compute-dominant (e.g., NVIDIA V100 to B200: 139 to 312 FLOP/Byte), the gap between compute and memory widens, so memory-focused designs like PAT become increasingly valuable; (2) model architecture. PAT yields large gains for common architectures that retain a KV cache (MHA, GQA), but benefits may shrink for architectures that compress or remove KV state (MLA (Liu et al., 2024), linear attention (Shen et al., 2021), MLKV (Zuhri et al., 2024)); and (3) prefix-sharing ratio in the batch. High concurrency with cross-request shared prefixes amplifies PAT ’s advantage, whereas small batches or workloads without shared prefixes limit the improvement (see Figure 9 1, 1920).
@@ 10. Related Work
KV Cache Related Optimization. Prior work reduced KV cache memory and fragmentation: FasterTransformer (NVIDIA, 2025d) provided an early static-batching implementation, Orca (Yu et al., 2022) improved utilization with continuous batching but relied on pre-allocated caches, which leads to memory fragmentation. vLLM introduced a paged KV cache to cut waste by virtual memory management (Kwon et al., 2023). SGLang’s prefix-reuse lets shared prefix KV blocks be reused across requests and is widely used in production (Zheng et al., 2024b). These approaches shrink KV memory costs but do not accelerate attention, which dominates latency as context and output lengths grow.
Attention Kernel Optimization. Fused, on-chip attention kernel implementations like FlashAttention and FlashInfer reduce global memory traffic by combining attention steps into a single kernel (Dao et al., 2022; Shah et al., 2024; Ye et al., 2025), but their query-centric (one-query-per-CTA) design cannot exploit workload-level prefix sharing. Subsequent works (Zhu et al., 2024; Ye et al., 2024b; Lin et al., 2024a; Zheng et al., 2024a) pack the single shared system prompt to reduce redundant memory accesses, but can not generalize to workloads with multi-level prefixes. Several recent works (Ye et al., 2024a; Yao et al., 2024; Juravsky et al., 2024) extend to multi-level prefixes but rely on simple packing strategies that overlook trade-offs between overhead and savings. (Pan et al., 2025) also addresses multi-level prefix optimization, but its compute-oriented packing cost model mismatches the memory-bound nature of decode attention, leading to suboptimal results. Furthermore, all these works use one-size-fits-all kernel design, leading to resource inefficiency. In contrast, PAT uses a memory-centric packing strategy with multi-tile kernels and multi-stream execution to reduce redundant memory accesses and improve resource utilization for dynamic workloads.
GPU Scheduling. GPU scheduling strategies further improve kernel efficiency by balancing compute and memory demands: (Kamath et al., 2025) fuses prefill and decode via virtual CTAs to expose resource complementarity, and (Zhu et al., 2025) groups SMs so different SMs specialize in specific tasks, raising utilization. These techniques are largely orthogonal to PAT and could be combined to further reduce execution inefficiencies. Work like (Song et al., 2024) shows that CTA-style scheduling on alternative accelerators (e.g., NPUs) can help eliminate the remaining resource bubbles in PAT ’s execution (§9). Here, we focus on GPU decode-attention optimizations and leave heterogeneous-scheduler integration to future work.

### v2

@@ 6. Multi Kernel Forward
Long KV Split. Multi-stream execution alone cannot eliminate execution bubbles in all cases, because some CTAs may have KV lengths that are orders of magnitude larger than others. We therefore adopt a KV-dimension splitting strategy similar to (pan2025fasttree). Specifically, we split any CTAs with KV length exceeding the mean KV length for all CTAs of the batch into equal parts to keep KV length under the mean value. This shortens the completion time of the last finishing CTAs and improves overall SM utilization of kernel execution.
@@ 7. Output Merge
We implement a lightweight merge kernel that uses online softmax (dao2022flashattention) to combine partial results at per-query granularity. Each CTA produces three per-query and per-head intermediates: a max score, a log-sum-exp accumulator, and a partial value-weighted sum. The merge kernel loads these intermediates from global memory, reduces them with online softmax, normalizes the accumulated value-weighted sum, concatenates all heads, and writes the final query output back to global memory. The small global read/write overhead for these intermediates is accounted for in the pack scheduler’s overhead analysis when deriving the end-to-end packing scheme (5.1).
@@ 8. Evaluation
In this section, we present the implementation of PAT (§8.1) and the experimental setup (§8.2). We then conduct a set of experiments to answer three key questions:
@@ 8.1. Implementation of PAT
We implement PAT as a full attention kernel for the decode stage with about 3k lines of Cutlass/CuTe (nvidia2025cutlassdocs; nvidia2025cutequickstart) and C++ code. The multi-tile kernel (§5.2), multi-stream forward (§6), and merge kernel (§7) are built with Cutlass/CuTe. The asynchronous pack scheduler (§5.1) and API wrappers are implemented in C++. To overlap the data transfers with computation, all data movement from the global memory to the shared memory uses the cp_async primitive, together with double buffering (nvidia2025cudaprogguide). We expose the kernel API to Python via pybind11, and integrate it into vLLM (kwon2023pagedattention) (v0.9.0) as an off-the-shelf plugin with about 1.2k lines of Python code. PAT treats vLLM’s paged KV cache as its substrate: KV entries are managed as fixed-size blocks in block tables, and the pack scheduler operates only on block IDs produced by vLLM. This design lets PAT reuse vLLM’s existing KV paging implementation. To enable PAT in vLLM, only an environment variable VLLM_ATTENTION_BACKEND=PAT is required.
@@ 8.2. Experiment Setup
Models and Testbed. For end-to-end online serving (§8.4), we use two representative open-source models: Qwen3-8B (yang2025qwen3) and Llama3-8B (meta2024llama3). All experiments run on the NVIDIA A100-80GB-SXM4 GPU. Hardware info is summarized in Table 1. Software environment is CUDA 12.4 and PyTorch 2.7.0.
Baselines. We compare against five attention implementations spanning query-centric and KV-centric designs:
  1. (1)
    FlashAttention (dao2022flashattention; dao2023flashattention2) (v2.6.2): query-centric; maps each query to a CTA with a fixed tile size config $(64,128)$.
  2. (2)
    FlashInfer (ye2025flashinfer) (v0.2.5): query-centric; improves SM load balance by dynamic CTA partitioning; decoding tile config $(16,128)$.
  3. (3)
    FastTree (pan2025fasttree): KV-centric; uses a compute-oriented cost model to pack and reduce repeated KV loads; two tile configs $(64,32)$ and $(16,32)$.
  4. (4)
    RelayAttention (zhu2024relayattention): KV-centric; packs single-level shared prefixes across requests into CTAs to cut redundant memory accesses and runs them with FlashAttention’s kernel.
  5. (5)
    RelayAttention++: our extension of RelayAttention to exploit vLLM-style KV-cache reuse; it stores shared KV blocks from non-first-level prefixes in the same physical space so redundant KV loads can benefit from L2 cache, further improving performance (§8.3).
Figure 11. End-to-end performance of PAT and the baselines under two models and two traces. Note that RelayAttention++ lacks support for multiple first-level prefixes, so its results on toolagent trace are unavailable. :
@@
Kernel-Performance Workloads (§8.3). To compare kernel performance under identical batch and prefix structures, we construct synthetic decode batches as input, following (pan2025fasttree). Each decode batch is specified by $B$ and $L$. $B$ defines the prefix-tree structure and the number of leaves (i.e., batch size). e.g., $B=[1,4,16]$ yields two shared-prefix levels with 1 and 4 nodes, and 16 leaves. $L$ gives KV lengths per level. e.g., $L=[128,256,1024]$ sets level-1 and level-2 shared prefixes to 128 and 256 tokens, with 1024 non-shared tokens per request. We vary $(B,L)$ combinations to reflect different shared-prefix structures and batch settings as in Figure 10. Besides, we choose four head configurations (#heads, #kv_heads) common in Llama (meta2024llama3), Qwen (yang2025qwen3), and Gemma (google2025gemma) models: $(64,8)$, $(32,8)$, $(16,8)$, and $(32,32)$. The head dimension and data type are set to the commonly used 128 and FP16.
@@
End-to-end workloads (§8.4 and §8.5). We evaluate PAT under online serving using vLLM (kwon2023pagedattention) (v0.9.0) on two real-world traces. (1) toolagent (qin2025mooncake): tool/agent workloads with task-specific system prompts (overall cache hit rate 59%). (2) conversation: combines the Meta-AI system instruction (total lengths 2517/2522 tokens for Llama3/Qwen3 tokenizers) with burstgpt prompts, following prior work (pan2025fasttree). We randomize language and country fields in the system instruction to create a three-level prefix whose lengths are 46, 348, 2123 (Llama3) or 45, 351, 2126 (Qwen3). We use only the first 30 minutes of each trace due to cost limit.
@@ 8.3. Kernel Performance
Overall results. Figure 10 reports normalized kernel performance (metric $\text{latency}^{-1}$, normalized to PAT) across decode-batch configurations. For configurations with shared prefixes ( 118), PAT achieves up to $21.5\times$, $11.7\times$, $3.2\times$, $11.9\times$ and $5.7\times$ speedups over FlashAttention, FlashInfer, FastTree, RelayAttention and RelayAttention++, respectively, across four attention-head settings. These gains arise from three factors: (1) a prefix-aware packing scheduler that cuts redundant global KV accesses (§5.1); (2) a multi-tile kernel combined with an online tile-size selector that adapts $(m,n)$ per CTA to better use bandwidth and on-chip memory (§5.2); and (3) multi-stream forward execution and long-KV splitting that mitigate execution bubbles caused by multi-tile kernels and KV-length dynamics (§6). Together, these designs make PAT consistently more efficient than the baselines.
Compared with query-centric kernels. Against query-centric FlashAttention and FlashInfer, PAT reduces attention latency by $67.8\%$ and $52.1\%$ on average under configurations with prefixes ( 118). The gap widens with larger batch size or longer shared prefixes (e.g., 45 and 67) because the one-query-per-CTA design of query-centric kernels forces repeated global KV loads that grow more costly in those cases. We also include configurations 19 and 20 in Figure 10, which remove shared prefixes and thus represent workloads without prefix reuse. In this case, PAT no longer reduces global memory accesses but still benefits from its multi-tile kernel and multi-stream forward. As a result, PAT achieves $1.6\%$ lower latency on average with fewer execution bubbles.
Compared with KV-centric kernels. FastTree, which also targets multi-level shared prefixes, remains the strongest baseline but is still $3.8\%$$68.9\%$ slower than PAT for two reasons: (1) its compute-oriented packing cost model is ill-suited to memory-bound decode attention (§2.3 and §3.2); and (2) its double-tile approach launches two kernels serially as in Figure 13(b), introducing execution bubbles. By contrast, PAT uses a memory-oriented packing strategy plus multi-stream forward to avoid these inefficiencies. RelayAttention++ cuts latency by $67.4\%$ versus RelayAttention, confirming that L2 plus KV-cache reuse reduces redundant global loads, yet RelayAttention++ is still $1.7\times$ slower than PAT, showing that L2 cache alone cannot fully eliminate redundant KV loads.
@@ 8.4. End-to-End Comparison
Overall trends. We compare PAT with three baselines across two models (Llama-3-8B and Qwen3-8B) and two real-world traces (conversation and toolagent) as in §8.2. Figure 11 shows the trend of mean TTFT, mean TPOT, and P99 TPOT as we vary the request rate. Specifically, as in subfigures b1 to b4, PAT reduces mean TPOT at the same request rate by 17.2–68.1% over RelayAttention++, 17.0–89.5% over FlashAttention, and 32.2–93.1% over FlashInfer. Furthermore, as shown by the first and the third rows of Figure 11, the TPOT reduction allows PAT to finish incoming requests faster at the same request rate. This yields 9.3–98.6%, 10.1–99.6%, and 22.5–99.8% lower TTFT than the three baselines.
@@ 8.5. Ablation Study
Figure 13. CTA execution pipeline on $SM_{0}-SM_{5}$ with two tile size configurations (collected by PTX (nvidia2023ptxisa)), where white space represents execution bubbles. (a) Multi-stream execution pipeline. (b) Serial execution pipeline.:
To evaluate the contribution of each design in PAT, we build the following ablation baselines: (1) PAT-compute, which adopts the cost model from FastTree (pan2025fasttree) for the packing scheduler; (2) PAT-naive, which simply packs each node in the tree-structured block table into a CTA; (3) PAT-fixed, which disables the multi-tile kernel of PAT and instead uses the fixed tile configuration $(64,128)$ as in FlashAttention (shah2024flashattention3); (4) PAT-serial, which disables the multi-stream forward mechanism of PAT and adopts the serial multi-kernel execution similar to FastTree (pan2025fasttree). We use the same synthetic traces as in §8.3 and adopt the attention head configuration of Llama-3-8B to compare PAT with ablation baselines.
@@
Importance of multi-tile kernel. As shown in Figure 12(a), enforcing a fixed tile size (PAT-fixed) increases attention latency by 39% compared to PAT. In the ablation workload, CTAs exhibit query sizes from 1 to 64 and KV lengths from 32 to 4096, which makes one-size-fits-all kernels highly inefficient. Prior kernels (shah2024flashattention3; ye2025flashinfer; zhu2024relayattention; yao2024deft) adopt such fixed designs, leading to padding overhead and execution bubbles. In contrast, PAT’s multi-tile kernel enables efficient adaptation to diverse CTA configurations.
@@
8.6. Overhead Analysis
The primary overhead introduced by PAT comes from the pack scheduler, which packs decode batches into CTAs at runtime. To mitigate its overhead, as described in §5.1, we design a lazy update mechanism that reduces the schedule’s triggering frequency without affecting correctness and overlaps its execution with the serving system’s preprocessing stage. To evaluate the overhead, we measure both the scheduling latency and the preprocessing latency under the toolagent and conversation traces at a request rate of 8 and 5 req/s. As shown in Figure 14, the average scheduling latency is lower than the preprocessing latency by 81.6% and 88.8%, respectively. Consistently, the P99 TPOT in Figure 11 c1- c4 shows 19.4–93.4% reduction compared with the baselines. Therefore, when running on an asynchronous CPU thread, the pack scheduler will not introduce additional end-to-end latency, demonstrating the effectiveness of the lazy update mechanism in PAT.
@@ 9. Discussion
Prospects and Limitations. PAT leverages cross-request prefix sharing to improve bandwidth utilization and cut redundant global memory accesses in memory-bound decode attention. Its effectiveness depends on three factors: (1) hardware compute-to-bandwidth ratio. As GPUs become more compute-dominant (e.g., NVIDIA V100 to B200: 139 to 312 FLOP/Byte), the gap between compute and memory widens, so memory-focused designs like PAT become increasingly valuable; (2) model architecture. PAT yields large gains for common architectures that retain a KV cache (MHA, GQA), but benefits may shrink for architectures that compress or remove KV state (MLA (liu2024deepseek), linear attention (shen2021efficient), MLKV (zuhri2024mlkv)); and (3) prefix-sharing ratio in the batch. High concurrency with cross-request shared prefixes amplifies PAT ’s advantage, whereas small batches or workloads without shared prefixes limit the improvement (see Figure 10 1, 1920).
@@ 10. Related Work
KV Cache Related Optimization. Prior work reduced KV cache memory and fragmentation: FasterTransformer (nvidia2025fastertransformer) provided an early static-batching implementation, Orca (yu2022orca) improved utilization with continuous batching but relied on pre-allocated caches, which leads to memory fragmentation. vLLM introduced a paged KV cache to cut waste by virtual memory management (kwon2023pagedattention). SGLang’s prefix-reuse lets shared prefix KV blocks be reused across requests and is widely used in production (zheng2024sglang). Recent systems (qin2024mooncake; lee2024infinigen) further extend KV capacity by offloading KV blocks between GPU and CPU or NVMe storage, which reduces on-GPU memory pressure at the cost of additional data movement. These approaches shrink KV memory costs but do not accelerate decode attention, which dominates latency as context and output lengths grow. PAT is orthogonal to these approaches because its pack scheduler only relies on logical KV block IDs, and the serving system transfers the required KV blocks to GPU memory before the attention kernel executes.
Attention Kernel Optimization. Fused, on-chip attention kernel implementations like FlashAttention and FlashInfer reduce global memory traffic by combining attention steps into a single kernel (dao2022flashattention; shah2024flashattention3; ye2025flashinfer), but their query-centric (one-query-per-CTA) design cannot exploit workload-level prefix sharing. Subsequent works (zhu2024relayattention; ye2024chunkattention; lin2024parrot; zheng2024batchllm) pack the single shared system prompt to reduce redundant memory accesses, but can not generalize to workloads with multi-level prefixes. Several recent works (zihao2024cascade; yao2024deft; juravsky2024hydragen) extend to multi-level prefixes but rely on simple packing strategies that overlook trade-offs between overhead and savings. (pan2025fasttree) also addresses multi-level prefix optimization, but its compute-oriented packing cost model mismatches the memory-bound nature of decode attention, leading to suboptimal results. Furthermore, all these works use one-size-fits-all kernel design, leading to resource inefficiency. In contrast, PAT uses a memory-centric packing strategy with multi-tile kernels and multi-stream execution to reduce redundant memory accesses and improve resource utilization for dynamic workloads.
GPU Scheduling. GPU scheduling strategies further improve kernel efficiency by balancing compute and memory demands: (kamath2025pod) fuses prefill and decode via virtual CTAs to expose resource complementarity, and (zhu2025nanoflow) groups SMs so different SMs specialize in specific tasks, raising utilization. These techniques are largely orthogonal to PAT and could be combined to further reduce execution inefficiencies. Work like (song2024tackling) shows that CTA-style scheduling on alternative accelerators (e.g., NPUs) can help eliminate the remaining resource bubbles in PAT ’s execution (§9). Here, we focus on GPU decode-attention optimizations and leave heterogeneous-scheduler integration to future work.